Apriori algorithm

Q2. In a retail store following sales transactions took place: 

 

Sales_ID Items S1 {O,W,C,N,Z} S2 {O,C,K,N,Z} S3 {K,N,W,Z,O} S4 {K,W,Z,O} S5 {C,K,N,Z} S6 {O,K,N} S7 {O,N,W,Z} S8 {K,N,Z} 

 

Find frequent itemsets and association rules using Apriori Algorithm. The minimum support is 0.4 (threshold) and minimum confidence is 0.9. 
Answer #1

Total cases \(=175\)

Probability that SUV price will increase when there is a decrease in tax rate

\(=\frac{30}{30+60} * \frac{60}{60+30}\)

\(=\frac{1}{3} * \frac{2}{3}=\frac{2}{9}\)

Probability that tax rate will increase when there is decrease in SUV price

\(=\frac{30}{85} *\left(\frac{55}{85}\right)\)

\(=\frac{1650}{7225}\)

\(=\frac{66}{288}\)

\(=\frac{11}{48}\)

Probabiity that tax rate decreases and suv price increases

\(=\frac{60}{90} * \frac{30}{90}=\frac{2}{9}\)

answered by: Kenkaindo